# Reward Inside the Model: A Lightweight Hidden-State Reward Model for LLM’s Best-of-N sampling

### v1

@@
Abstract
High-quality reward models are crucial for unlocking the reasoning potential of large language models (LLMs), with best-of-N voting demonstrating significant performance gains. However, current reward models, which typically operate on the textual output of LLMs, are computationally expensive and parameter-heavy, limiting their real-world applications. We introduce the Efficient Linear Hidden State Reward (ELHSR) model - a novel, highly parameter-efficient approach that leverages the rich information embedded in LLM hidden states to address these issues. ELHSR systematically outperform baselines with less than 0.005% of the parameters of baselines, requiring only a few samples for training. ELHSR also achieves orders-of-magnitude efficiency improvement with significantly less time and fewer FLOPs per sample than baseline reward models. Moreover, ELHSR exhibits robust performance even when trained only on logits, extending its applicability to some closed-source LLMs. In addition, ELHSR can also be combined with traditional reward models to achieve additional performance gains.
1 Introduction
The recent “reasoning era” (Chen et al., 2025) in large language models (LLMs) has witnessed a remarkable surge in their ability to tackle complex reasoning tasks. Models such as GPT-o3-mini (OpenAI, 2024), DeepSeek-R1 (Guo et al., 2025), Grok 3 (xAI, 2025), Claude 3.7 Sonnet (Anthropic, 2025) and Gemini 2.5 (Google, 2025) have demonstrated remarkably impressive reasoning performance. Moreover, recent studies have highlighted the considerable potential of using reward model guidance to enhance the reasoning capabilities of LLMs, typically through best-of-N sampling due to its simplicity and effectiveness (Snell et al., 2025; Brown et al., 2024).
However, current methods generally rely on predicting reward scores based on the LLM’s generated text using independently trained neural networks - often with substantial parameter counts. However, these approaches incur significant overhead, as training and deploying large reward models consume substantial computational resources and require large amount of data during both the reward model training phase and the subsequent inference stage (Guo et al., 2025; Namgoong et al., 2024; Zhang et al., 2024). This exclusive reliance on text-based information constrains these models, necessitating large training sets and complex architectures to adequately capture the full spectrum of LLM errors.
However, during the reasoning process, the LLM possesses an inherent ‘inner knowledge’ reflecting its current certainty. Existing research demonstrates that the LLM’s internal states during reasoning contain high levels of information about answer confidence, often captured through linear representations (Zhang et al., 2025; Feucht et al., 2024; Xie et al., 2024). Our preliminary experiment in Section 3.1 also supports this observation by showing that hidden states can encode a linear representation for reasoning’s correctness with high accuracy.
Figure 1: An illustration comparing traditional reward model and ELHSR.
While prior work has utilized hidden states to improve LLM’s reasoning performance, they often focus on the final answer token (Burns et al., 2023; Azaria and Mitchell, 2023; Xie et al., 2024) or on specific answer tokens within each reasoning step (Zhang et al., 2025). This approach has limitations: relying on a single token’s hidden state may not capture the full context of the reasoning process, and defining clear reasoning steps and identifying their corresponding answer tokens can be subjective and challenging. In addition, they do not provide direct reward signals, and their performance gains are limited compared to large-scale reward models. Therefore, we propose the Efficient Linear Hidden State Reward (ELHSR) model (illustrated in Figure 1). ELHSR learns a linear gating and a linear reward projection per token. The final reward is computed as a weighted average of these token-level rewards, scaled by the gating values.
Our experiments demonstrate that ELHSR systematically outperforms baseline reward models while using fewer than 0.005% their parameters. Moreover, ELHSR model is data-efficient and can be trained using only a few instances. Critically, ELHSR achieves orders-of-magnitude improvements of efficiency in average time and FLOPs per sample compared to baselines.
In summary, this work makes three main contributions:
  • We introduce ELHSR, a novel token-level reward model based on LLM hidden states. ELHSR achieves competitive reasoning performance with orders-of-magnitude less parameters and computational cost.
  • ELHSR demonstrates exceptional data efficiency and scalability, performing well with limited data, demonstrating performance upgrade with training-time scaling and test-time scaling, generalizing to logit-only training for certain closed-source LLMs, and can be combined with conventional reward models to yield further performance improvements.
  • We analyze ELHSR’s mechanisms by providing concrete examples, illustrating the way that it scores the critical components of reasoning paths.
2 Problem Statement
@@
A common approach to enhance LLM reasoning is best-of-N sampling (Snell et al., 2025). Let $\mathcal{V}$ be the token vocabulary and let $D$ be a distribution over problem-answer pairs $(x,y)\in\mathcal{V}^{n}\times\mathcal{Y}$, where $y$ is the hidden reference solution for question $x$. Conditioned on $x$, the base LLM $\pi$ draws $N$ reasoning paths independently (i.i.d. draws from $\pi(\cdot\mid x)$), producing $r_{i}=(r_{i,1},\dots,r_{i,m_{i}})\in\mathcal{V}^{m_{i}},\quad i=1,\dots,N,$ where $m_{i}$ is the random length of the $i$-th path. We write $\mathcal{V}^{*}=\bigcup_{\ell=0}^{\infty}\mathcal{V}^{\ell}$. Simultaneously, $\pi$ emits hidden states $h_{i}=(h_{i,1},\dots,h_{i,m_{i}})\;\in\;(\mathbb{R}^{L\times d})^{m_{i}},$ where $L$ is the number of transformer layers, $d$ is the per-layer hidden-state dimension, and each $h_{i,j}\in\mathbb{R}^{L\times d}$ is the stacked hidden state after emitting $r_{i,j}$. Define the correctness indicator $F:\mathcal{V}^{*}\times\mathcal{Y}\to\{0,1\},\quad\text{s.t.}\quad F(r,y)=1\iff r\text{ yields the reference answer }y.$ We write $(\mathbb{R}^{L\times d})^{*}=\bigcup_{k=0}^{\infty}(\mathbb{R}^{L\times d})^{k}$. A parametric reward model $R_{\theta}:(\mathbb{R}^{L\times d})^{*}\to\mathbb{R}$ assigns a scalar score to each $h_{i}$ (hence for each $r_{i}$). Note that neither $\pi$ nor $R_{\theta}$ sees $y$. At test time we select $i^{*}=\arg\max_{1\leq i\leq N}R_{\theta}(h_{i})$ and return $r_{i^{*}}$. Our goal is to choose $\theta$ to maximize the probability that $r_{i^{*}}$ is correct (expectation of $F(r_{i^{*}},y)$) under $D$:
$\max_{\theta}\;\mathbb{E}_{(x,y)\sim D}\;\mathbb{E}_{(r_{1},\dots,r_{N})\sim\pi(\cdot\mid x)}\bigl{[}F\bigl{(}r_{\arg\max_{i}R_{\theta}(h_{i})},y\bigr{)}\bigr{]}.$ (1)
3 Methods
3.1 Motivation
@@
Figure 2: Cross-validation accuracy of the PCA+LDA pipeline for predicting correctness at each layer of Llama-3.1-8B. The results demonstrate that hidden states contain information about reasoning correctness with linear representation.
@@
We first present a preliminary experiment demonstrating the potential for enhancing reasoning using hidden states. Recent work has shown that LLM hidden states contain information about the correctness probability of a given reasoning path, often exhibiting linear representations (Xie et al., 2024; Zhang et al., 2025; Feucht et al., 2024). To verify this, we designed a simple pipeline to predict the correctness of reasoning steps based on the LLM’s hidden states.
@@
Our pipeline extracts hidden states $h^{\ell}_{t}$ for each token $t$ at each layer $l$ from the Llama-3.1-8B-Instruct model (Grattafiori et al., 2024) during reasoning on the MATH dataset (Hendrycks et al., 2021), averages the hidden states across all tokens within each reasoning path to obtain a single vector representation $\bar{h}^{\ell}$ for each layer $\ell$, then reduces $\bar{h}^{\ell}$ for each $\ell$ to 50 dimensions using Principal Component Analysis (PCA) (Wold et al., 1987), and classifies the reasoning as correct or incorrect using Linear Discriminant Analysis (LDA) (Balakrishnama and Ganapathiraju, 1998). The pipeline is linear to verify that the hidden states can encode a linear representation for confidence of the answer. For more details, please refer to Appendix C.2.
@@
3.2 Our Approach - ELHSR Model
Therefore, building upon the observation that LLM internal representations exhibit linear properties, we designed Efficient Linear Hidden State Reward (ELHSR) to be a parameter-efficient, linear model.
Our approach, ELHSR, learns a token-level reward function directly from the concatenated hidden states. ELHSR is designed to be highly parameter-efficient.
@@ 3.2 Our Approach - ELHSR Model
$\begin{pmatrix}\tilde{g}_{t}\\ r_{t}\end{pmatrix}=W_{ELHSR}h_{t}+b_{ELHSR}$ (2)
where $W_{ELHSR}\in\mathbb{R}^{2\times Ld}$ is the weight matrix, $b_{ELHSR}\in\mathbb{R}^{2}$ is the bias vector, $\tilde{g}_{t}$ is the pre-activation gating value, and $r_{t}$ is the token-level reward. The gating value is then passed through a sigmoid function (Han and Moraga, 1995):
$g_{t}=\sigma(\tilde{g}_{t})=\frac{1}{1+e^{-\tilde{g}_{t}}}$ (3)
@@
$R=\frac{\sum_{t=1}^{T}(g_{t}\cdot r_{t})}{\max\left(\sum_{t=1}^{T}g_{t},\epsilon\right)}$ (4)
@@
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\cdot\log(\sigma(R(h_{i})))+(1-y_{i})\cdot\log(1-\sigma(R(h_{i})))\right]$ (5)
@@
Although binary cross-entropy with logits loss generally yields good results, our experiments detailed in Appendix E.1 demonstrate that DPO (Rafailov et al., 2023) and hinge loss (Rosasco et al., 2004) can also perform well, sometimes even surpassing cross-entropy. However, given that cross-entropy performs best on average, we adopt it as our primary loss function in the main text. A comparison study among several loss functions is provided in Appendix E.1
4 Experiments
4.1 Experimental Setup
We evaluate our method on the MATH (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), and AQuA_RAT (Ling et al., 2017) datasets. The models used for generation are Llama-3.2-3B-Instruct (Meta, 2024), Llama-3.1-8B-Instruct (Grattafiori et al., 2024), and Ministral-8B-Instruct (Mistral AI, 2024). The generation temperature is 1.0 with top_p to be 0.9. We use 6000 instances (each with 8 reasoning paths) for training and 500 instances for evaluation. When training ELHSR, we use early stopping (Yao et al., 2007) based on validation set performance to mitigate overfitting. For more details on the generation process, training and evaluation, please refer to Appendix C.3, C.4 and C.5.
We compare our method against four recent open-source reward models: EurusRM-7B (Yuan et al., 2025), Skywork-Reward-Llama-3.1-8B-v0.2 (Liu et al., 2024), Starling-7B (Zhu et al., 2024) and UltraRM-13B (Cui et al., 2024). These baselines are very parameter heavy and trained on extensive datasets, while achieving leading performance among open-source reward models. For links and brief descriptions of baselines, please refer to Appendix A. For parameter count and training data size of baselines compared against ELHSR, see Table 2 for details.
To provide a more comprehensive comparison, we also evaluated a fine-tuned version of EurusRM-7B, detailed in Appendix C.6, specifically fine-tuned on the relevant datasets and model-generated solutions used in our experiments. This is despite EurusRM-7B already being pre-trained on MATH and GSM8K, which gives it a potential advantage.
4.2 Accuracy
Table 1: Different methods’ best-of-N sampling performance on MATH, GSM8K, and AQuA_RAT datasets. Results represent accuracy (percentages omitted). Our method outperforms baselines. Bold: best performance; Underline: second-best performance; @k: indicates the number of reasoning paths used in best-of-N sampling.

### v3

# Mining Intrinsic Rewards from LLM Hidden States for Efficient Best-of-N Sampling

@@
Abstract.
Best-of-N sampling is a powerful method for improving Large Language Model (LLM) performance, but it is often limited by its dependence on massive, text-based reward models. These models are not only computationally expensive but also data-hungry, requiring extensive labeled datasets for training. This creates a significant data challenge, as they overlook a rich, readily available data source: the LLM’s own internal hidden states. To address this data and efficiency gap, we introduce SWIFT (Simple Weighted Intrinsic Feedback Technique), a novel and lightweight method that learns a reward function directly from the rich information embedded in LLM hidden states. Operating at the token embedding level, SWIFT employs simple linear layers to effectively distinguish between preferred and dispreferred generations, eliminating the need for computationally intensive text-based modeling. Extensive experiments on standard benchmarks show that SWIFT outperforms existing baselines (12.7% higher accuracy than EurusRM-7B on MATH dataset) while using less than 0.005% of their parameters. Its robust scalability, compatibility with certain closed-source models via logit access, and ability to combine with traditional reward models for additional performance highlight SWIFT’s practical value and contribution to more efficient data-driven LLM post-training. Our code is available at https://github.com/aster2024/SWIFT.
Large Language Models, Knowledge Discovery, Reward Modeling
journalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 09–13, 2026; Jeju Island, Republic of Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’26), August 09–13, 2026, Jeju Island, Republic of Koreadoi: 10.1145/3770854.3780302isbn: 979-8-4007-2258-5/2026/08ccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Natural language processingccs: Computing methodologies Learning from implicit feedbackccs: Computing methodologies Learning latent representations
Resource Availability:
The source code of this paper has been made publicly available at https://doi.org/10.5281/zenodo.18140886.
1. Introduction
Figure 1. An illustration comparing traditional reward model and SWIFT.
An illustration comparing traditional reward model and SWIFT.
The proliferation of Large Language Models (LLMs) has unlocked unprecedented capabilities in various domains, from complex reasoning to user assistance (Liu et al., 2023; Zhang et al., 2023; Samel et al., 2023). To further refine and enhance these capabilities after initial pretraining, various post-training techniques are employed, including methods that scale computation at test-time to improve performance on difficult tasks (Liu et al., 2025). A key one of these test-time scaling strategies is Best-of-N sampling, in which a reward model selects the best response from a set of candidates (Snell et al., 2025). However, the effectiveness of this approach is significantly constrained by the reward models themselves. Current methods rely on large, independently trained neural networks that predict rewards based on generated text. These models are not only computationally expensive but also data-hungry, requiring massive datasets and incurring substantial overhead during both training and inference (Guo et al., 2025; Namgoong et al., 2024; Zhang et al., 2024). This exclusive reliance on textual data necessitates complex architectures to capture subtle errors, posing a major barrier to efficient LLM enhancement.
We argue that this paradigm overlooks a rich, readily available data source: the task-performing LLM’s own internal hidden states. During generation, an LLM naturally emits internal signals that reflect its confidence of its answer and the quality of its reasoning. Seminal work has shown that this “internal knowledge” is not only information-rich but is often captured in a linearly separable representations within the model’s hidden states (Zhang et al., 2025; Xie et al., 2024). Our own preliminary experiments support this finding (Section 4.1), revealing a clear linear relationship between hidden states and reasoning correctness. This property is significant, as it points to a path for highly efficient reward modeling: rather than collecting ever-larger text datasets to train ever-larger external models, we can mine the intrinsic signals already present within the LLM itself.
However, leveraging this insight introduces a fundamental architectural challenge. Traditional reward models are designed to process discrete text tokens, creating a mismatch with the continuous, high-dimensional vectors of hidden states. Addressing this mismatch requires a new architecture tailored specifically to extract meaningful signals from these hidden representations. While prior work has explored hidden states to improve reasoning, such methods typically focus on single, final-answer tokens or rely on subjective definitions of reasoning steps (Burns et al., 2023; Azaria and Mitchell, 2023). As a result, they offer limited performance gains. They also do not produce a direct reward signal suitable for Best-of-N sampling.
To address this challenge, we propose the Simple Weighted Intrinsic Feedback Technique (SWIFT), a novel architecture specifically designed to learn from hidden states at the token level (Figure 1). For each token, SWIFT learns both a linear gating value and a linear reward projection. The final reward is computed as a weighted average of these token-level rewards, offering a holistic, fine-grained assessment of the entire generation. Our experiments show that SWIFT outperforms baseline reward models (demonstrated 12.7% higher accuracy in average on MATH dataset than EurusRM-7B, a state-of-the-art open-source reward model on math and code domain (Yuan et al., 2025)) while using fewer than 0.005% of their parameters. It is exceptionally data-efficient, can be trained with only a handful of examples, and achieves orders-of-magnitude improvements in computational efficiency (FLOPs) compared to existing baselines.
Our code is available at https://github.com/aster2024/SWIFT, and the project website is on https://aster2024.github.io/swift-website/. In summary, our main contributions are:
  • We introduce SWIFT, a token-level reward model that directly mines LLM hidden states. We demonstrate SWIFT’s orthogonal design can augment traditional reward models to achieve further performance gains.
  • We extend SWIFT to operate with logit-only access, demonstrating its practical applicability to a broader range of models, including certain closed-source LLMs.
  • Our experiments demonstrate that SWIFT is highly data-efficient, achieving strong performance with minimal training data. Furthermore, it exhibits robust scalability with performance improvement from both training-time and test-time scaling.
2. Related Work
Reward Modeling for Language Models’ Reasoning
The rapid advancement of Large Language Models (LLMs) in complex tasks, such as answering user’s questions (Shen et al., 2023; Mo et al., 2023) and reasoning (Zhou et al., 2024; Xu et al., 2024), has highlighted the pivotal role of reward modeling in both training and post-training inference (Cui et al., 2024; Lambert et al., 2025). Current reward models typically build upon an LLM backbone, with substantial parameter count and require significant amount of data for training, which introduces high computational costs (Guo et al., 2025; Namgoong et al., 2024). Although prior work has proposed various approaches to mitigate such issue (Chan et al., 2024; Fisch et al., 2024), achieving a satisfactory balance between efficiency and performance remains a challenge. This work introduces a novel perspective by leveraging internal representations to achieve a significant improvement in efficiency with satisfactory performance.
The rapid development of LLMs has spurred considerable research into their internal representations (Ferrando et al., 2024; Li et al., 2023). Hidden states of LLMs have been effectively utilized in diverse fields, including safety alignment (Zou et al., 2023; Barez et al., 2025), faithfulness and hallucination detection (Burns et al., 2023), knowledge editing (Meng et al., 2022; Yu et al., 2024), and knowledge distillation or transfer (Dasgupta and Cohn, 2025; Wu et al., 2024). By contrast, we effectively leverage the hidden states in intermediate layers of LLMs as a direct signal for reward modeling in reasoning tasks.
3. Problem Statement
@@
A common approach to enhance LLM reasoning is best-of-N sampling (Snell et al., 2025). Let $\mathcal{V}$ be the token vocabulary and let $D$ be a distribution over problem-answer pairs $(x,y)\in\mathcal{V}^{n}\times\mathcal{Y}$, where $y$ is the hidden reference solution for question $x$. Conditioned on $x$, the base LLM $\pi$ draws $N$ reasoning paths independently (i.i.d. draws from $\pi(\cdot\mid x)$), producing $r_{i}=(r_{i,1},\dots,r_{i,m_{i}})\in\mathcal{V}^{m_{i}},\quad i=1,\dots,N,$ where $m_{i}$ is the random length of the $i$-th path. We write $\mathcal{V}^{*}=\bigcup_{\ell=0}^{\infty}\mathcal{V}^{\ell}$. Simultaneously, $\pi$ emits hidden states $h_{i}=(h_{i,1},\dots,h_{i,m_{i}})\;\in\;(\mathbb{R}^{L\times d})^{m_{i}},$ where $L$ is the number of transformer layers, $d$ is the per-layer hidden-state dimension, and each $h_{i,j}\in\mathbb{R}^{L\times d}$ is the stacked hidden state after emitting $r_{i,j}$. Define the correctness indicator $F:\mathcal{V}^{*}\times\mathcal{Y}\to\{0,1\},\text{s.t.}\,F(r,y)=1\iff r\text{ yields reference answer }y.$ We write $(\mathbb{R}^{L\times d})^{*}=\bigcup_{k=0}^{\infty}(\mathbb{R}^{L\times d})^{k}$. A parametric reward model $R_{\theta}:(\mathbb{R}^{L\times d})^{*}\to\mathbb{R}$ assigns a scalar score to each $h_{i}$ (hence for each $r_{i}$). Note that neither $\pi$ nor $R_{\theta}$ sees $y$. At test time we select $i^{*}=\arg\max_{1\leq i\leq N}R_{\theta}(h_{i})$ and return $r_{i^{*}}$. Our goal is to choose $\theta$ to maximize the probability that $r_{i^{*}}$ is correct (expectation of $F(r_{i^{*}},y)$) under $D$:
(1) $\max_{\theta}\;\mathbb{E}_{(x,y)\sim D}\;\mathbb{E}_{(r_{1},\dots,r_{N})\sim\pi(\cdot\mid x)}\bigl[F\bigl(r_{\arg\max_{i}R_{\theta}(h_{i})},y\bigr)\bigr].$
4. Methods
4.1. Motivation
@@
Figure 2. Cross-validation accuracy of the PCA+LDA pipeline for predicting correctness at each layer of Llama-3.1-8B. The results demonstrate that hidden states contain information about reasoning correctness with linear representation.
Cross-validation accuracy of the PCA+LDA pipeline for predicting correctness at each layer of Llama-3.1-8B.
@@
We first present a preliminary experiment demonstrating the potential for enhancing reasoning using hidden states. Recent work has shown that LLM hidden states contain information about the correctness probability of a given reasoning path, often exhibiting linear representations (Zhang et al., 2025; Xie et al., 2024). To verify this, we designed a simple pipeline to predict the correctness of reasoning steps based on the LLM’s hidden states.
@@
Our pipeline extracts hidden states $h^{\ell}_{t}$ for each token $t$ at each layer $l$ from the Llama-3.1-8B-Instruct model (Grattafiori et al., 2024) during reasoning on the MATH dataset (Hendrycks et al., 2021), averages the hidden states across all tokens within each reasoning path to obtain a single vector representation $\bar{h}^{\ell}$ for each layer $\ell$, then reduces $\bar{h}^{\ell}$ for each $\ell$ to 50 dimensions using Principal Component Analysis (PCA) (Wold et al., 1987), and classifies the reasoning as correct or incorrect using Linear Discriminant Analysis (LDA) (Balakrishnama and Ganapathiraju, 1998). The pipeline is linear to verify that the hidden states can encode a linear representation for confidence of the answer. For more details, please refer to Appendix D.2.
@@
4.2. Our Approach - SWIFT
Therefore, building upon the observation that LLM internal representations exhibit linear properties, we designed SWIFT (Simple Weighted Intrinsic Feedback Technique) model to be a parameter-efficient, linear model, which learns a token-level reward function directly from the concatenated hidden states.
@@ 4.2. Our Approach - SWIFT
(2) $\begin{pmatrix}\tilde{g}_{t}\\ r_{t}\end{pmatrix}=W_{SWIFT}h_{t}+b_{SWIFT}$
where $W_{SWIFT}\in\mathbb{R}^{2\times Ld}$ is the weight matrix, $b_{SWIFT}\in\mathbb{R}^{2}$ is the bias vector, $\tilde{g}_{t}$ is the pre-activation gating value, and $r_{t}$ is the token-level reward. The gating value is then passed through a sigmoid function (Han and Moraga, 1995):
(3) $g_{t}=\sigma(\tilde{g}_{t})=\frac{1}{1+e^{-\tilde{g}_{t}}}$
@@
(4) $R=\frac{\sum_{t=1}^{T}(g_{t}\cdot r_{t})}{\max\left(\sum_{t=1}^{T}g_{t},\epsilon\right)}$
@@
(5) $\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\cdot\log(\sigma(R(h_{i})))+(1-y_{i})\cdot\log(1-\sigma(R(h_{i})))\right]$
@@
Although binary cross-entropy with logits loss generally yields good results, our experiments detailed in Appendix E.1 demonstrate that DPO (Rafailov et al., 2023) and hinge loss (Rosasco et al., 2004) can also perform well, sometimes even surpassing cross-entropy. However, given that cross-entropy performs best on average, we adopt it as our primary loss function in the main text. A comparison study among several loss functions is provided in Appendix E.1.
5. Experiments
5.1. Experimental Setup
Our experiments are conducted across mathematical reasoning, code comprehension, commonsense abilities and symbolic reasoning tasks. We evaluate our method on the MATH (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), AQuA_RAT (Ling et al., 2017), Imbue Code Comprehension (Imbue Team, 2024), HellaSwag (Zellers et al., 2019) and CoinFlip (Wei et al., 2022) datasets. The models used for generation are Llama-3.2-3B-Instruct (Meta, 2024), Llama-3.1-8B-Instruct (Grattafiori et al., 2024), and Ministral-8B-Instruct (Mistral AI, 2024). The generation temperature is 1.0 with top_p to be 0.9. We use 6000 instances (each with 8 reasoning paths) for training and 500 instances for evaluation. When training SWIFT, we use early stopping (Yao et al., 2007) based on validation set performance to mitigate overfitting. For more details on the generation process, training and evaluation, please refer to Appendix D.3, D.4 and D.5.
We compare our method against six recent open-source reward models: EurusRM-7B (Yuan et al., 2025), Skywork-Reward-Llama-3.1-8B-v0.2 (Liu et al., 2024), Starling-7B (Zhu et al., 2024), UltraRM-13B (Cui et al., 2024), RLHFlow-8B-Deepseek (Xiong et al., 2024), Math-Shepherd-7B (Wang et al., 2024). These baselines are very parameter heavy and trained on extensive datasets, while achieving leading performance among open-source reward models. These six baseline reward models have all been explicitly trained on data related to reasoning and code domains. Their enhanced performance in these specific areas is well-documented and validated in their respective original publications. It is particularly noteworthy that the development of Eurus-7B, Ultra-13B, RLHFlow-8B-Deepseek, and Math-Shepherd-7B involved training and evaluation on the MATH and GSM8K datasets. This corresponds directly with the datasets used in our work, confirming the relevance of these models as strong baselines.
For links and brief descriptions of baselines, please refer to Appendix B. For parameter count and training data size of baselines compared against SWIFT, see Table 3 for details.
To provide a more comprehensive comparison, we also evaluated a fine-tuned version of EurusRM-7B in Table 12, with details illustrated in Appendix D.6, specifically fine-tuned on the relevant datasets and model-generated solutions used in our experiments, while using the same loss function as in the original publication to ensure consistency. This is despite EurusRM-7B already being pre-trained on MATH and GSM8K, which gives it a potential advantage.
5.2. Accuracy
Table 1. Different methods’ best-of-N sampling performance on mathematics domain, including MATH, GSM8K, and AQuA_RAT datasets. Results represent accuracy (percentages omitted). Bold: best performance; Underline: second-best performance; @k: indicates the number of reasoning paths used in best-of-N sampling.
Different methods’ best-of-N sampling performance on mathematics domain, including MATH, GSM8K, and AQuA_RAT datasets.